Ricardo deposits $4500 in a trust fund that pays 15% interest at the end
of each year.
a. Write a recursive rule that describes how the amount of money in
Ricardo’s trust fund increases over time.

Answer by ikleyn(52847)

.

The recursive rule is   A(n+1) = 1.15*A(n),   n= 1, 2, 3, . . . ,   from year to year.


Solved.